# Probing false vacuum decay and bubble nucleation in a Rydberg atom array

### v1

@@ Abstract
In quantum field theory (QFT), the “vacuum” is not just empty space but the lowest-energy state of a quantum field. If the energy landscape has multiple local minima, the local ground states are the false vacuum (FV) which can tunnel towards the global ground state (true vacuum, TV). This process exhibits signature akin to classical supercooled gas transitions and many-body tunneling in discrete quantum systems. Here, we study the FV decay and bubble nucleation in a Rydberg atom ring. The long-range van-der-Waals interactions and individual-site addressability allow us to explore physics beyond the standard Ising model. We observe that the FV decay rate decreases exponentially with the inverse of the symmetry-breaking field, directly mirroring QFT predictions. Moreover, we demonstrate that even minor deviations from the ideal metastable state can cause a stark departure from this universal scaling law. Extending beyond short-time decay dynamics, we also examine resonant bubble nucleation, a feature distinctive to systems with discrete energy spectra. Our findings and methods open avenues for future studies of many-body tunneling in higher dimensions or more complex geometries.
Introduction. — Quantum tunneling in many-body systems [RevModPhys.93.025005, 2016_Anders_DirectQuantumTunneling, Boixo2016, RevModPhys.85.367, RevModPhys.80.885, 2025_YinChao_MetastableState, Meng2015] offers dynamics intrinsically richer than those of single particles or collective phases described by a single order parameter. A distinct feature of such tunneling is the nucleation and growth of locally transformed regions, or “bubbles” [Langer1967CondensationPoint, Kobzarev1974Bubbles] within the initial metastable state [2021_Lagnese_FVDinSpinChains, 2021_Sinha_QuantizedBubbleNucleation, Lencs_2022_tricritical, 2024_Darbha_AFM]. The study of this phenomenon has recently become a topical pursuit for quantum simulation platforms of increasing scale and capability [2024_Zenesini_SuperFluid, 2025_Vodeb_DWave, 2024_ZhenShengYuan_OpticalLattice, 2025_Monroe_irontrap].
The false vacuum decay (FVD) in lattice systems is among the most intensely studied quantum many-body tunneling phenomena. Its dynamics are often analogized to the instanton description in quantum field theory (QFT) [1977_Coleman_Instanton, PhysRevD.16.1762, devoto2022false, 1990_Farhi_CreatUniverse], where spontaneous symmetry breaking destabilizes a ground state. The decay of such metastable states is ubiquitous, appearing in diverse physical systems  [KRAMERS1940, Langer1969, Binder1987, RevModPhys1990, Debenedetti2001, Klix2010, Macieszczak2016, Defenu_2021] and underscoring its broad relevance. Numerous experimental schemes are proposed to realize FVD, utilizing platforms such as ultracold atoms  [2017Fialko_BEC, 2018Braden_BEC, Billam_2019, Ng_2021], trapped ions  [2013Hauke_trappedions, 2016Yang_trappedions], and superconducting circuits [2021Abel_superconducting]. Recent experimental progress includes observation of bubble emergence [2024_Zenesini_SuperFluid, 2025_Monroe_irontrap], the study of bubble hopping and interaction dynamics in weak coupling [2025_Vodeb_DWave], and the identification of excitation modes via Fourier spectroscopy of quench dynamics [2024_ZhenShengYuan_OpticalLattice].
@@
Rydberg atom arrays have recently emerged as a powerful and versatile platform for simulating quantum many-body dynamics [2010_Saffman_RMP, 2022_Saffman_QuantumComputer, 2021_Lukin_256atom, 2023_ChenCheng_CSB, 2025_KKN_CriticalPhenomena]. These systems, featuring individually trapped and detected cold atoms in optical tweezers with tunable interactions between Rydberg states, can be utilized to simulate diverse spin models [2025_Yue_Disorder, 2025_Xinhui_OTOC, 2025_QiaoMu_TJmodel]. Furthermore, the ability to individually address and manipulate atoms with lasers enables the exploration of even richer quantum phases and non-equilibrium phenomena [2024_Browaeys_LocalControl, 2025_XuPeng_FiberArray, 2025_LinLi_LatticeGaugeTheory].
@@
In this work, we simulate FVD in an antiferromagnetic (AFM) Ising ring using a Rydberg atom array. By applying a site-dependent longitudinal field, we explicitly break the $\mathbb{Z}_{2}$ symmetry of the two Néel ground states, casting one as a tunable false vacuum (FV) and the other as the true vacuum (TV). We first probe the short-time dynamics of a Néel state after a quench, observing the characteristic exponential decay of AFM order. We then study the dynamics of a more refined metastable state, the pre-quenched ground (PQG) state, and find it exhibits a much stabler decay. Supported by numerical simulations, we show that the PQG decay rate follows the theoretically predicted exponential suppression with the inverse symmetry-breaking field. This behavior holds even for a generalized Ising model incorporating both global and staggered potentials, as well as long-range interactions. Finally, we explore the long-time dynamics to study resonant bubble nucleation, a process characteristic of systems with a discrete energy spectrum.
@@
Figure 1: Simulating false vacuum decay with a programmable Rydberg atom array. (a) Schematic of the experimental setup. Neutral atoms are trapped in a ring geometry and illuminated globally by the 420-nm and 1013-nm lasers, coupling the ground state to a high-lying Rydberg state (70S). A set of far-detuned, site-selective addressing lasers (pink beams) illuminates every other atom, generating a staggered detuning. A magnetic field $B$ parallel to the 420-nm laser defines the quantization axis. (b) Energy landscape for false vacuum decay in an antiferromagnetic Ising model. The staggered longitudinal field breaks the degeneracy of the two Néel-ordered ground states, creating a metastable false vacuum and a stable true vacuum. Quantum tunneling from the false vacuum towards the true vacuum proceeds via the nucleation of true vacuum bubbles (illustrated as gray domains of flipped spins). The dashed lines denote static energies ($\Omega=0$) of representative spin configurations along the unfolded ring.
As illustrated in Fig. 1(a), our system is a ring-shaped array consisting of $N$ (with $N$ even) uniformly spaced single $^{87}\rm{Rb}$ atoms, with the pseudo-spin-$\frac{1}{2}$ encoded in the ground state $|\!\downarrow\rangle\equiv|5S_{1/2},F=2,m_{F}=2\rangle$ and the Rydberg state $|\!\uparrow\rangle\equiv|70S_{1/2},m_{J}=1/2\rangle$. The spin states are coupled via a two-photon transition (420 nm and 1013 nm), through the intermediate state $|6P_{3/2},F=3,m_{F}=3\rangle$, with an effective Rabi frequency $\Omega$. The Hamiltonian governing the system is
@@
where $\hat{n}_{j}=|\!\uparrow\rangle_{j}\!\langle\uparrow\!|$ ($\hat{g}_{j}=|\!\downarrow\rangle_{j}\!\langle\downarrow\!|$) is the Rydberg (ground) state occupation operator, $\hat{\sigma}^{x}_{j}=|\!\uparrow\rangle_{j}\!\langle\downarrow\!|+\text{h.c.}$ is the spin-flip operator. The term $V_{i,j}\propto|i-j|^{-6}$ describes the van der Waals interaction, which is antiferromagnetic ($V_{i,j}>0$) and dominated by the nearest-neighbor coupling $V\equiv V_{i,i+1}$. The global detuning $\Delta_{g}$ and staggered detuning $(-1)^{j}\Delta_{l}$ are engineered by adjusting the two-photon detuning of 420-nm and 1013-nm lasers, in conjunction with the AC Stark shifts of site-selective addressing laser beams [Fig. 1(a)]. Retaining only the nearest neighbor interactions and setting $\Delta_{g}=V$, our antiferromagnetic Hamiltonian maps directly to the well studied ferromagnetic quantum Ising model with both transverse and longitudinal fields: $H_{f}=-\sum_{j=1}^{N}\left[\sigma^{z}_{j}\sigma^{z}_{j+1}+h_{x}\hat{\sigma}^{x}_{j}+h_{z}\hat{\sigma}^{z}_{j}\right]$, with $h_{x}=2\Omega/V$ and $h_{z}=2\Delta_{l}/V$. Crucially, our platform is not limited to this mapping; by freely tuning $\Delta_{g}\neq V$, we can explore a generalized Ising model and physics beyond the standard Ising model.
@@
In the absence of the staggered field ($\Delta_{l}=0$) and in the limit $\Omega\to 0$, the Hamiltonian $\hat{H}$ possesses a $\mathbb{Z}_{2}$ symmetry and features two degenerate Néel-ordered ground states. The staggered field $\Delta_{l}$ acts as the symmetry-breaking field, lifting this degeneracy (Fig. 1(b)). For $\Delta_{l}>0$, the state $|\text{Néel}\rangle=|\!\downarrow\uparrow\dots\downarrow\uparrow\rangle$ becomes the high-energy metastable state (the FV), while $|\text{Néel}^{\prime}\rangle=|\!\uparrow\downarrow\dots\uparrow\downarrow\rangle$ becomes the low-energy TV.
@@
The decay from the FV proceeds via the nucleation of TV bubbles, as sketched in Fig. 1(b). The smallest bubble can form in two ways: (i) flipping an $|\!\uparrow\rangle$ spin into $|\!\downarrow\rangle$, acquiring an energy of $\Delta E_{\uparrow\to\downarrow}=\Delta_{g}-\Delta_{l}$; or (ii) flipping a $|\!\downarrow\rangle$ spin into $|\uparrow\rangle$, causing an energy penalty of $\Delta E_{\downarrow\to\uparrow}=2V-\Delta_{g}-\Delta_{l}$. These energy costs become degenerate, $\Delta E=V-\Delta_{l}$, in the standard Ising limit $\Delta_{g}=V$. For this standard model, a TV bubble of length $L$ has an energy cost of $V-L\Delta_{l}$. Resonant nucleation occurs when this cost vanishes, $L\Delta_{l}\approx V$.
@@
FV decay of the Néel state in a generalized Ising model — We begin by studying FVD in the generalized Ising regime with $\Delta_{g}=0.8V$. We prepare the initial $|\text{Néel}\rangle$ state (the FV) by applying AC Stark shifts to every other atom, followed by a Landau-Zener sweep [1932_Landau] to excite the addressed atoms to the Rydberg state (see Supplemental Material (SM) Sec. I [SM]). The decay dynamics are initiated by a quench of $\Omega$, $\Delta_{g}$ and $\Delta_{l}$, after which we track the decay of the AFM order parameter,
$\langle\hat{M}_{\mathrm{AFM}}\rangle=\frac{1}{N}\sum_{j}(-1)^{j}\langle\hat{\sigma}_{j}^{z}\rangle.$ (2)
This parameter, which is 1 for the FV and -1 for the TV, is rescaled as in [2021_Lagnese_FVDinSpinChains]
@@
Figure 2: Dynamics and scaling of the false vacuum decay. Main panel: Measured time evolution of the rescaled antiferromagnetic (AFM) order for an $N=16$ atom ring, plotted on a logarithmic scale for different $V/\Delta_{l}$ ratios. Experimental data (solid squares) are averaged over 300 realizations. Solid lines are exponential fits $M_{\text{AFM}}^{\text{~res}}(t)\propto e^{-\gamma t}$ to the data within time windows $t\in[0.02,0.18]~\rm{\mu}s$, used to extract the decay rate $\gamma$. Dashed lines are extensions of these fits. Inset: Extracted decay rate $\gamma$, normalized by $\Omega$, as a function of the inverse local staggered field $V/\Delta_{l}$. Experimental results for $N=16$ (empty diamonds) and $N=24$ (solid hexagons) are shown with statistical error bars. The gray line is the decay rate extracted from numerical simulations for $N=16$ without considering any experimental errors.
@@
The inset displays the extracted $\gamma$ (empty diamonds) alongside the results attained from numerical simulations for $N=16$ without considering any experimental errors (gray line). We repeated the measurement for a larger $N=24$ ring (solid hexagons) and found that the decay rates collapse onto the same curve. This size-independence suggests a universal scaling behavior for a sufficiently long chain. Such behavior is consistent with the local nature of the decay dynamics: in the short-time regime, the Lieb-Robinson bound [1972_Lieb_FiniteGroupVelocity, 2012_Bloch_LightConeCorrelation, 2015_Eisert_OutOfEquilibrium] restricts the decay process to a local light-cone, making the overall decay rate insensitive to the total system size $N$. Our numerical simulations corroborate this, showing that $M_{\text{AFM}}^{\text{res}}(t)$ and $\gamma$ converge for $N>12$ (except for very small $\Delta_{l}$).
@@
Besides, Figure 2 shows that the FV decay rate diminishes as $\Delta_{l}$ decreases. This mechanism can be understood heuristically with the aid of Fig. 1(b): the false vacuum’s energy becomes resonant with product states containing a length-$L$ bubble when the bulk-energy gain $L\Delta_{l}$ compensates the surface-energy cost $V$. This defines a critical bubble size $L_{c}\sim V/\Delta_{l}$. Since generating such a bubble requires a collective tunneling of $L_{c}$ spins, the nucleation rate $\gamma$ is expected to be exponentially suppressed with $L_{c}$. In deed, instanton theory describing the FV decay of scalar quantum field [1977_Coleman_Instanton, devoto2022false, PhysRevD.97.105002] and non-perturbative analytical study of the ferromagnetic Ising model [1999_Rutkevich_PerturbativeIsingModel] both predict an exponential suppression $\gamma\propto\exp(-\lambda V/\Delta_{l})$, where $\lambda$ collects the remaining dependencies. For an Ising chain, $\lambda$ depends only on the transverse field $\Omega$ [2021_Lagnese_FVDinSpinChains]. However, our results in Fig. 2 show a significant deviation from this well-established exponential prediction, particularly in the small-$\Delta_{l}$ regime.
@@
Figure 3: Comparison of FV decay of an initial Néel state (blue squares) versus the pre-quench ground (PQG) state (orange pentagons). (a) Measured evolution of the AFM order starting from the Néel state versus PQG state, for $\Delta_{l}/V=0$ (left) and $\Delta_{l}/V=0.08$ (right). Solid lines connecting data points are guides to the eye. Dashed lines are numerical simulations that include decoherence at the bench-marked level. (b) Comparison of decay rates, extracted from numerical simulation, for the Néel and PQG states as a function of $V/\Delta_{l}$. The left panel shows the result for $\Delta_{g}/V=1$, and the solid gray line is the theoretical prediction based on Eq. (6) of Ref. [2021_Lagnese_FVDinSpinChains] without any fitting parameters. The right panel shows the result for $\Delta_{g}/V=0.8$, and the dashed gray line represents a linear fit for the PQG data.
@@
Different initial state. — To understand the discrepancy observed in Fig. 2, we investigate the decay dynamics from a different state: the pre-quench ground (PQG) state. It is the entangled ground state when $\Delta_{l}\to 0^{-}$ (i.e., in the opposite symmetry-broken sector) and therefore serves as a more faithful representation of the metastable false vacuum. Experimentally, we prepare the $|\rm PQG\rangle$ state by first initializing the system in the $|\text{Néel}\rangle$ state and then adiabatically ramping up $\Omega$ while holding $\Delta_{l}=0$ (see SM Sec. I [SM]). We then quench $\Delta_{l}$ to its target value and observe the decay process.
@@
Figure 3(a) directly compares the measured decay of the PQG state (orange pentagons) against the Néel state (blue squares) for both $\Delta_{l}/V=0$ and $\Delta_{l}/V=0.08$. The PQG state exhibits a significantly slower decay. This behavior is qualitatively confirmed by numerical simulations that incorporate decoherence (dashed lines in Fig. 3(a)) (see SM Sec. II [SM]).
@@
To gain deeper insight into these different dynamics, we numerically simulate the post-quench evolution of both states for various $\Delta_{l}$. We extract the decay rates $\gamma$ by fitting $M_{\text{AFM}}^{\text{res}}(t)$ over a $\Delta_{l}$-distinctive time window  [^1] We extract $\gamma$ from numerical simulations by fitting $\ln[M_{\rm{AFM}}^{\rm{~res}}(t)]$ within a dynamically determined window $t\in[t_{\rm{start}},t_{\rm{end}}]$, where $t_{\rm{start}}=0.034+0.0058V/\Delta_{l}~(\rm{\mu s})$, $t_{\rm{end}}=0.11+0.034V/\Delta_{l}~(\rm{\mu s)}$ exhibiting an exponential decay (see SM Sec. III [SM]). The results are shown in Fig. 3(b) for two scenarios: the standard Ising model ($\Delta_{g}=V$, nearest-neighbor interactions only), for which an analytical expression for $\gamma$ is known for small longitudinal fields [1999_Rutkevich_PerturbativeIsingModel, 2021_Lagnese_FVDinSpinChains], and (ii) the generalized Ising model ($\Delta_{g}=0.8V$, long-range van der Waals interactions), matching our experimental parameters from Fig. 2.
In the left panel of Fig. 3(b), we compare the decay rate extracted from numerical simulations with theoretical predictions using Eq. (6) of Ref. [2021_Lagnese_FVDinSpinChains] (solid gray line). Remarkably, the decay rates for the PQG state show excellent agreement with the theory over four orders of magnitude, with no fitting parameters. In contrast, for the Néel state, it becomes difficult to identify a clear exponential decay regime when $V/\Delta_{l}\geq 10$ (small $\Delta_{l}$). This is because the Néel state exhibits overwhelming fluctuations that obscure the weak FVD signal (see SM Sec. III  [SM]). These fluctuations arise from the excess energy of the Néel state relative to the PQG state, which cannot be dissipated in an isolated system. Nevertheless, when using the same time windows (as the PQG’s) to forcibly fit $M_{\rm{AFM}}^{\rm{~res}}(t)$, we find that the resulted $\gamma$ for the Néel state rapidly deviate from the theoretical prediction as $\Delta_{l}$ decreases. These striking distinctions, present despite the high overlap between the two states, can be understood heuristically: for small $\Delta_{l}$, the PQG state is, by construction, nearly an eigenstate of the post-quench Hamiltonian, ensuring its metastability. In contrast, the Néel state corresponds to the groundstate when $\Delta_{l}\to+\infty$, leaving it a broader superposition of eigenstates of the post-quench Hamiltonian.
@@
Scenario (ii) ($\Delta_{g}=0.8V$ with long-range interactions) maps to a generalized Ising model with both global and staggered longitudinal fields, for which no analytical expression for $\gamma$ is known. In addition, long-range van-der-Waals interaction between atoms are considered. Nevertheless, our simulations (Fig. 3(b), right) show that the same dichotomy persists: the PQG state’s decay rate still follows a clear exponential scaling, $\gamma\propto\exp(-\lambda V/\Delta_{l})$ (dashed gray fit), albeit with a different slope $\lambda$, while the Néel state does not. This discovery suggests that the exponential suppression of $\gamma$ with symmetry-breaking field strength $\Delta_{l}$ is a universal feature of FVD in generalized Ising model, provided the system is prepared in a proper metastable state. The universality of this law in general 1D systems and the precise role of long-range interactions in FVD dynamics are open questions for future exploration.
[…]

### v2

@@ Abstract
In quantum field theory (QFT), the “vacuum” is not just empty space but the lowest-energy state of a quantum field. If the energy landscape has multiple local minima, the local ground states are the false vacuum (FV) which can tunnel towards the global ground state (true vacuum, TV). This process exhibits signature akin to classical supercooled gas transitions and many-body tunneling in discrete quantum systems. Here, we study the FV decay and bubble nucleation in a Rydberg atom ring. The $1/r^{6}$ van-der-Waals interactions and individual-site addressability allow us to explore physics beyond the standard Ising model. We observe that the FV decay rate decreases exponentially with the inverse of the symmetry-breaking field, directly mirroring QFT predictions. Moreover, we demonstrate that even minor deviations from the ideal metastable state can cause a stark departure from this universal scaling law. Extending beyond short-time decay dynamics, we also examine resonant bubble nucleation, a feature distinctive to systems with discrete energy spectra. Our findings and methods open avenues for future studies of many-body tunneling in higher dimensions or more complex geometries.
Introduction. — Quantum tunneling in many-body systems [1, 2, 3, 4, 5] exhibits dynamics intrinsically richer than those of single particles or collective phases described by a single order parameter. A paradigmatic example is false vacuum decay (FVD) in quantum field theory (QFT). This process involves the tunneling of a metastable vacuum towards a lower-energy state, which is typically described by instanton solutions [6, 7, 8, 9, 10]. A distinct feature of such tunneling is the nucleation and growth of “bubbles”—locally transformed regions—within the metastable state [11, 12, 13, 14]. Such decay processes are ubiquitous across diverse physical systems, spanning from classical to quantum, continuous to discrete, open to closed [15, 16, 17, 18]. Potential realizations have been explored in ultracold atoms [19, 20, 21], trapped ions [22, 23], and superconducting circuits [24]. Recent experimental progress includes observation of bubble emergence [25, 26], bubble hopping and interaction dynamics in weak coupling [27], and excitation modes via Fourier spectroscopy of quench dynamics [28].
@@
Rydberg atom arrays have recently emerged as a powerful and versatile platform for simulating quantum many-body dynamics [29, 30, 31, 32, 33]. These systems, featuring individually trapped and detected atoms with tunable interactions between Rydberg states, can be utilized to simulate diverse spin models [34, 35, 36]. Furthermore, the ability to individually address and manipulate atoms with lasers enables the exploration of even richer quantum phases and non-equilibrium phenomena [37, 38, 39].
@@
In this work, we simulate FVD in an antiferromagnetic (AFM) Ising ring using a programmable Rydberg atom array. By applying a site-dependent longitudinal field, we lift the $\mathbb{Z}_{2}$ degeneracy of the two Néel ground states, casting one as a false vacuum (FV) and the other as the true vacuum (TV). We first probe the short-time dynamics of a Néel state after a quench, observing the characteristic exponential decay of AFM order. We then study the dynamics of a more refined metastable state, the pre-quench ground (PQG) state. We show that the PQG decay rate is exponentially suppressed by the inverse symmetry-breaking field, a feature that persists down to the weak-field regime and contrasts sharply with the Néel state. This behavior holds even for a generalized Ising model incorporating both global and staggered potentials, as well as $1/r^{6}$ van-der-Waals interactions. Finally, we explore the long-time dynamics to study resonant bubble nucleation, a process characteristic of systems with a discrete energy spectrum.
@@
Figure 1: Simulating false vacuum decay with a programmable Rydberg atom array. (a) Schematic of the experimental setup. Neutral atoms are trapped in a ring geometry and illuminated globally by the 420-nm and 1013-nm lasers, coupling the ground state to a high-lying Rydberg state (70S). A set of far-detuned, site-selective addressing lasers (pink beams) illuminates every other atom, generating a staggered detuning. A magnetic field $B$ parallel to the 420-nm laser defines the quantization axis. (b) Energy landscape for false vacuum decay in an antiferromagnetic Ising model. The staggered longitudinal field breaks the degeneracy of the two Néel-ordered ground states, creating FV and TV states. Quantum tunneling from the FV towards the TV proceeds via the nucleation of TV bubbles (illustrated as gray domains of flipped spins). The dashed lines denote static energies ($\Omega=0$) of representative spin configurations along the unfolded ring.
As illustrated in Fig. 1(a), our system is a ring-shaped array consisting of $N$ (with $N$ even) uniformly spaced $^{87}\rm{Rb}$ atoms, with the pseudo-spin-$\frac{1}{2}$ encoded in the ground state $|\!\downarrow\rangle\equiv|5S_{1/2},F=2,m_{F}=2\rangle$ and the Rydberg state $|\!\uparrow\rangle\equiv|70S_{1/2},m_{J}=1/2\rangle$. The spin states are coupled via a two-photon transition (420-nm and 1013-nm), through the intermediate state $|6P_{3/2},F=3,m_{F}=3\rangle$, with an effective Rabi frequency $\Omega$. The Hamiltonian governing the system is
@@
where $\hat{n}_{j}=|\!\uparrow\rangle_{j}\!\langle\uparrow\!|$ ($\hat{g}_{j}=|\!\downarrow\rangle_{j}\!\langle\downarrow\!|$) is the Rydberg (ground) state occupation operator, $\hat{\sigma}^{x}_{j}=|\!\uparrow\rangle_{j}\!\langle\downarrow\!|+\text{h.c.}$ is the spin-flip operator. The term $V_{i,j}\propto|i-j|^{-6}$ describes the van-der-Waals interaction, which is antiferromagnetic ($V_{i,j}>0$) and dominated by the nearest-neighbor coupling $V\equiv V_{i,i+1}$. The global detuning $\Delta_{g}$ and staggered detuning $(-1)^{j}\Delta_{l}$ are engineered by adjusting the two-photon detuning of 420-nm and 1013-nm lasers, in conjunction with the AC Stark shifts of site-selective addressing laser beams [Fig. 1(a)]. Retaining only the nearest neighbor interactions and setting $\Delta_{g}=V$, the antiferromagnetic Hamiltonian (1) maps directly to the well studied ferromagnetic quantum Ising model with both transverse and longitudinal fields: $\hat{H}_{f}/\hbar=-\sum_{j=1}^{N}\left[\hat{\sigma}^{z}_{j}\hat{\sigma}^{z}_{j+1}+h_{x}\hat{\sigma}^{x}_{j}+h_{z}\hat{\sigma}^{z}_{j}\right]$, with $h_{x}=2\Omega/V$ and $h_{z}=2\Delta_{l}/V$. Crucially, by setting $\Delta_{g}\neq V$, we can explore a generalized Ising model and physics beyond the standard Ising model with this platform.
@@
In the absence of the staggered field ($\Delta_{l}=0$) and in the limit $\Omega\to 0$, the Hamiltonian $\hat{H}$ possesses a $\mathbb{Z}_{2}$ symmetry and features two degenerate Néel-ordered ground states. The staggered field $\Delta_{l}$ acts as the symmetry-breaking field, lifting this degeneracy (Fig. 1(b)). For $\Delta_{l}>0$, the state $|\text{Néel}\rangle=|\!\downarrow\uparrow\dots\downarrow\uparrow\rangle$ becomes the high-energy metastable state (the FV), while $|\text{Néel}^{\prime}\rangle=|\!\uparrow\downarrow\dots\uparrow\downarrow\rangle$ becomes the low-energy TV  [^1] In the presence of a transverse field $\Omega$, the classical Néel states are not exact eigenstates; they serve as approximations of the vacuum states..
@@
The decay from the FV proceeds via the nucleation of TV bubbles, as sketched in Fig. 1(b). The smallest bubble can form in two ways: (i) flipping an $|\!\uparrow\rangle$ spin into $|\!\downarrow\rangle$, acquiring an energy of $\Delta E_{\uparrow\to\downarrow}=\Delta_{g}-\Delta_{l}$; or (ii) flipping a $|\!\downarrow\rangle$ spin into $|\uparrow\rangle$, causing an energy penalty of $\Delta E_{\downarrow\to\uparrow}=2V-\Delta_{g}-\Delta_{l}$. These energy costs become degenerate, $\Delta E=V-\Delta_{l}$, in the standard Ising limit $\Delta_{g}=V$.
@@
FV decay of the Néel state in a generalized Ising model — We begin by studying FVD in the generalized Ising regime with $\Delta_{g}=0.8V$. We prepare the initial $|\text{Néel}\rangle$ state (the FV) by applying AC Stark shifts to every other atom, followed by a Landau-Zener sweep [41] to excite the addressed atoms to the Rydberg state (see Supplemental Material (SM) Sec. I [42]). The decay dynamics are initiated by a quench of $\Omega$, $\Delta_{g}$ and $\Delta_{l}$, after which we track the decay of the AFM order parameter,
$\langle\hat{M}_{\mathrm{AFM}}\rangle=\frac{1}{N}\sum_{j=1}^{N}(-1)^{j}\langle\hat{\sigma}_{j}^{z}\rangle.$ (2)
This parameter, which is 1 for the FV and -1 for the TV, is rescaled as in reference [12]
@@
Figure 2: Dynamics and scaling of the false vacuum decay. Main panel: Measured time evolution of the rescaled antiferromagnetic (AFM) order for an $N=16$ atom ring, plotted on a logarithmic scale for different $V/\Delta_{l}$ ratios. Experimental data (solid squares) are averaged over 300 realizations. Solid lines are exponential fits $M_{\text{AFM}}^{\text{~res}}(t)\propto e^{-\gamma t}$ to the data within time windows $t\in[0.02,0.18]~\rm{\mu}s$, used to extract the decay rate $\gamma$. Dashed lines are extensions of these fits. Inset: Extracted decay rate $\gamma$, normalized by $\Omega$, as a function of the inverse local staggered field $V/\Delta_{l}$. Experimental data for $N=16$ (empty diamonds) and $N=24$ (solid hexagons) are presented on a semi-logarithmic scale with statistical error bars. The gray line is the decay rate extracted from numerical simulations for $N=16$ without considering any experimental errors.
@@
The inset displays the extracted $\gamma$ (empty diamonds) alongside the results attained from numerical simulations for $N=16$ without considering any experimental errors (gray line). We repeated the measurement for a larger $N=24$ ring (solid hexagons) and found that the decay rates collapse onto the same curve. This size-independence suggests a universal scaling behavior for a sufficiently long chain. Such behavior is consistent with the local nature of the decay dynamics: in the short-time regime, the Lieb-Robinson bound [54, 55, 56] restricts the decay process to a local light-cone, making the overall decay rate insensitive to the total system size $N$. Our numerical simulations corroborate this, showing that $M_{\text{AFM}}^{\text{res}}(t)$ and $\gamma$ converge for $N>12$ (except for very small $\Delta_{l}$).
@@
Besides, Figure 2 shows that the FV decay rate diminishes as $\Delta_{l}$ decreases. This mechanism can be understood heuristically with the aid of Fig. 1(b): the false vacuum’s energy becomes resonant with product states containing a length-$L$ bubble when the bulk-energy gain $L\Delta_{l}$ compensates the surface-energy cost $V$. This defines a resonant bubble size $L_{r}\sim V/\Delta_{l}$. Since generating such a bubble requires a collective tunneling of $L_{r}$ spins, the nucleation rate $\gamma$ is expected to be exponentially suppressed with $L_{r}$. Indeed, instanton theory describing the FV decay of scalar quantum field [6, 10] and non-perturbative analytical study of the ferromagnetic Ising model [57] both predict an exponential suppression $\gamma\propto\exp(-\lambda V/\Delta_{l})$, where $\lambda$ collects the remaining dependencies. For a standard Ising chain, $\lambda$ depends only on the transverse field $\Omega$ [12]. Intriguingly, our data in Fig. 2 exhibit a notable deviation from this exponential prediction, failing to collapse onto a straight line in the semi-log plot, especially at small $\Delta_{l}$.
@@
Figure 3: Comparison of FV decay of an initial Néel state (blue squares) versus the pre-quench ground (PQG) state (orange pentagons). (a) Measured evolution of the AFM order starting from the Néel state versus PQG state, for $\Delta_{l}/V=0$ (left) and $\Delta_{l}/V=0.08$ (right). Dashed lines are numerical simulations that include decoherence at the bench-marked level. (b) Comparison of decay rates, extracted from numerical simulation, for the Néel and PQG states as a function of $V/\Delta_{l}$. The left panel shows the result for $\Delta_{g}/V=1$ (nearest neighbor only), and the solid gray line is the theoretical prediction based on Ref. [12]. The right panel shows the result for $\Delta_{g}/V=0.8$ (full van-der-Waals interactions), and the dashed gray line represents a linear fit for the PQG data.
@@
Different initial states. — To understand the discrepancy observed in Fig. 2, we investigate the decay dynamics from a different state: the pre-quench ground (PQG) state. It is the entangled ground state when $\Delta_{l}\to 0^{-}$ (i.e., in the opposite symmetry-broken sector) and therefore serves as a more faithful representation of the metastable false vacuum. Experimentally, we prepare the $|\rm PQG\rangle$ state by first initializing the system in the $|\text{Néel}\rangle$ state and then adiabatically ramping up $\Omega$ while holding $\Delta_{l}=0$ (see SM Sec. I [42]). We then quench $\Delta_{l}$ to its target value and observe the decay process.
@@
Figure 3(a) directly compares the measured decay of the PQG state (orange pentagons) against the Néel state (blue squares) for both $\Delta_{l}/V=0$ and $\Delta_{l}/V=0.08$. The PQG state exhibits a significantly slower decay. This behavior is qualitatively confirmed by numerical simulations that incorporate decoherence (dashed lines in Fig. 3(a)) (see SM Sec. II [42]). The different early-time dynamics of the two initial states can also be revealed by analytical analysis via a Baker-Campbell-Hausdorff (BCH) expansion (see SM Sec. IV [42]).
@@
To gain deeper insight into these different dynamics, we numerically simulate the post-quench evolution of both states for various $\Delta_{l}$. We extract the decay rates $\gamma$ by fitting $M_{\text{AFM}}^{\text{res}}(t)$ over a $\Delta_{l}$-distinctive time window  [^2] We extract $\gamma$ from numerical simulations by fitting $\ln[M_{\rm{AFM}}^{\rm{~res}}(t)]$ within a dynamically determined window $t\in[t_{\rm{start}},t_{\rm{end}}]$, where $t_{\rm{start}}=0.034+0.0058V/\Delta_{l}~(\rm{\mu s})$, $t_{\rm{end}}=0.11+0.034V/\Delta_{l}~(\rm{\mu s)}$ exhibiting an exponential decay (see SM Sec. III [42]). The results are shown in Fig. 3(b) for two scenarios: (i) the standard Ising model ($\Delta_{g}=V$, nearest-neighbor interactions only), for which an analytical expression for $\gamma$ is known for small longitudinal fields [57, 12], and (ii) the generalized Ising model ($\Delta_{g}=0.8V$, $1/r^{6}$ van-der-Waals interactions beyond the nearest-neighbor), matching our experimental parameters from Fig. 2.
@@
In the left panel of Fig. 3(b), we compare the decay rate extracted from numerical simulations with theoretical predictions from Eq. (6) of Ref. [12] (solid gray line, without any fitting parameters). Remarkably, the decay rates for the PQG state show excellent agreement with the theory over four orders of magnitude, with no fitting parameters. In contrast, for the Néel state, it becomes difficult to identify a clear exponential decay regime when $V/\Delta_{l}\geq 10$. This is because the weak exponential decay signal is masked by stronger high-frequency oscillations (see SM Sec. III [42]). Nevertheless, using the same time windows (as the PQG’s) to fit $M_{\rm{AFM}}^{\rm{~res}}(t)$, we find that the resulted $\gamma$ for the Néel state rapidly deviate from the theoretical prediction as $\Delta_{l}$ decreases. These distinctions are striking given the high overlap between the two states, but can be understood heuristically: for small $\Delta_{l}$, the PQG state is, by construction, nearly an eigenstate of the post-quench Hamiltonian, ensuring its metastability; in contrast, the Néel state, which corresponds to the ground state when $\Delta_{l}\to+\infty$, represents a broad superposition of eigenstates of the post-quench Hamiltonian.
@@
Scenario (ii) maps to a generalized Ising model with both global and staggered longitudinal fields, for which no analytical expression for $\gamma$ is known. Nevertheless, our simulations (Fig. 3(b), right) show that the same dichotomy persists: the PQG state’s decay rate still follows a clear exponential scaling, $\gamma\propto\exp(-\lambda V/\Delta_{l})$ (dashed gray fit), albeit with a different slope $\lambda$, while the Néel state does not. This discovery suggests that the exponential suppression of $\gamma$ with symmetry-breaking field strength $\Delta_{l}$ is a universal feature of FVD in generalized Ising model, provided the system is prepared in a proper metastable state. Recent theoretical works have proved the rigorous theoretical bounds of this decay law in general short-range interacting systems [5]. However, the precise role of long-tail interactions in FVD dynamics, as present in our Rydberg platform, remains an open question.